OpenStudy (anonymous):

8 times the square of a nonzero number is equal to 96 times the number. What is the number? A) 12 B) 24 C) 96 D) 4 @jdoe0001

OpenStudy (anonymous):

I believe it's 12, correct?

OpenStudy (jdoe0001):

12? hmm have you covered simplifying linear equations yet?

OpenStudy (anonymous):

Well if you plug it into the equation that it gives us, "8 * n^2 = 96(n)" if you replace n with 12, and do 8 * (12)^2 then that equals 1,152 So does 96(12) = 1,152.. So that's how I predicted the answer is 12.

OpenStudy (jdoe0001):

well... close , you want to "isolate" the "x" on the left-side you'd divide both sides by 8 then you take square root to both sides

OpenStudy (anonymous):

Oh okay.

OpenStudy (jdoe0001):

\(\bf 8\cdot x^2=96\implies x^2=\cfrac{96}{8}\implies x^2=12\\ \quad \\ \textit{taking }\sqrt{\qquad } \textit{ root to both sides}\\ \quad \\ \sqrt{x^2}=\sqrt{12}\implies x=\sqrt{12}\qquad {\color{blue}{ 12\implies 2\cdot 2\cdot 3\implies 2^2\cdot 3}}\\ \quad \\ x=\sqrt{12}\implies x=\sqrt{2^2\cdot 3}\implies x=\sqrt[{\color{red}{ 2}}]{2^{\color{red}{ 2}}\cdot 3}\implies x=2\sqrt{3}\)

OpenStudy (anonymous):

So the answer is 12 right?

OpenStudy (jdoe0001):

the nonzero number above would be "x"

OpenStudy (jdoe0001):

hhmmm wait is not one of your choices... so maybe is not asking for that

OpenStudy (anonymous):

That's why I said the answer IS 12, CORRECT?

OpenStudy (anonymous):

Yup it was 12, thanks.

OpenStudy (jdoe0001):

... I see... I missed one variable... but, yes is 12